# Mesh-Informed Reduced Order Models for Aneurysm Rupture Risk Prediction

### v2

@@ Abstract
The complexity of the cardiovascular system needs to be accurately reproduced in order to promptly acknowledge health conditions; to this aim, advanced multifidelity and multiphysics numerical models are crucial. On one side, Full Order Models (FOMs) deliver accurate hemodynamic assessments, but their high computational demands hinder their real-time clinical application. In contrast, ROMs provide more efficient yet accurate solutions, essential for personalized healthcare and timely clinical decision-making. In this work, we explore the application of computational fluid dynamics (CFD) in cardiovascular medicine by integrating FOMs with ROMs for predicting the risk of aortic aneurysm growth and rupture. Wall Shear Stress (WSS) and the Oscillatory Shear Index (OSI), sampled at different growth stages of the thoracic aortic aneurysm, are predicted by means of Graph Neural Networks (GNNs). GNNs exploit the natural graph structure of the mesh obtained by the Finite Volume (FV) discretization, taking into account the spatial local information, regardless of the dimension of the input graph. Our experimental validation framework yields promising results, confirming our method as a valid alternative that overcomes the curse of dimensionality.
@@ 1 Introduction
Thoracic aortic aneurysms (TAAs) represent a severe form of aneurysm that manifests as unrepairable bulges in the thoracic artery. These are particularly prevalent among the elderly, necessitating careful and regular monitoring of their size, especially in the case of small TAAs [3]. The progression and potential rupture of aneurysms are critical concerns, and their management often involves balancing the risks of surgical intervention against the potential for catastrophic vascular failure.
@@ 2.1 FOM
This section introduces the mathematical model used to describe blood flow, specifically the incompressible Navier-Stokes equations. This includes adopted boundary conditions and discretization methods.
@@ 2.1.1 Mathematical Model
The blood motion is determined within a time-independent domain $\Omega$ over a time interval $[t_{0},t_{F}]$. The flow is governed by the following equations:
$\begin{array}[]{ll}\rho\partial_{t}\boldsymbol{u}+\rho\nabla\cdot(\boldsymbol{u}\otimes\boldsymbol{u})-\nabla\cdot\boldsymbol{\sigma}=0&\text{ in }\Omega\times\left(t_{0},t_{F}\right],\end{array}$ (1)
@@
where $\rho=1060\,\text{kg/m}^{3}$ is the blood density, $\mathbf{u}$ represents velocity vector, and $\mathbf{\sigma}$ denotes the Cauchy stress tensor. Eq. 1 expresses the conservation of momentum, while Eq. 2 ensures the mass conservation. The blood is modeled as a non-Newtonian fluid, where $\mathbf{\sigma}$ is defined as:
@@
where $\mathbf{D}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{\mathrm{T}}\right)$, $p$ represents the pressure, $\mu$ is the dynamic viscosity and $\dot{\gamma}$ is the shear rate:
@@
where $\boldsymbol{\tau}_{0}$ and $\mu_{\infty}$ are 0.005 and 0.0035, respectively.
@@
where $\mathbf{n}$ is the normal vector, and $\boldsymbol{\tau}_{w}$ denotes the wall shear stress tensor.
@@
For the sake of simplicity, in what follows we denote by WSS the point-wise norm of $\mathbf{WSS}$.
@@ 2.1.3 Numerical discretization
To evaluate the independence of CFD results from mesh size, four sets of meshes containing $1.2\times 10^{5}$, $2.4\times 10^{5}$, $3.3\times 10^{5}$ and $4.5\times 10^{5}$ tetrahedral elements were created. These meshes correspond to varying element sizes on the aneurysm wall, specifically $0.125$ mm, $0.15$ mm, $0.2$ mm, and $0.3$ mm, respectively. Temporal variations of the area-averaged WSS on the aneurysm wall for all mesh sizes are presented in Fig. 3. This figure shows minimal differences between the two finest meshes, leading to the selection of the third mesh for subsequent simulations.
@@
Three cardiac cycles are required to achieve fully stationary results. Additionally, to ensure convergence, initial values of zero for pressure and velocity were applied, making the early cycles inaccurate.
@@ 3.1 Experimental setup
    Connectivity: The connectivity of the surface of the TAA is extracted from the Finite Volume Discretization, leading to the reference graph.The number of nodes and the connectivity depends on the specific size; values are reported in Table 1 ;
@@
    Node features: for each timestamp and for each TAA size, we label the graph nodes $v$ with a feature $\mathbf{x}_{v}=[x,y,z,t,v(t),p_{\text{abd}}(t),p_{\text{SA}}(t)]$, where: $x,y,z$ are the coordinates of each node (that varies according to the TAA size), $t$ is the times step, $v(t)$ is the inlet flow rate, $p_{\text{abd}}(t),p_{\text{SA}}(t)$ are the abdominal and SA outlet pressure.
@@
    the first step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 100%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 75%. This task focus on the interpolation capabilities of the model over shear indices at different growth stages of the TAA.
@@
    the second step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 75%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 100%. This task focus on the extrapolation capabilities of the model over shear indices at different growth stages of the TAA. This task is clinically more relevant as we would like to predict the WSS on future (and therefore larger) volume sizes of the TAA.
@@
From a qualitative point of view, we compare how GNNs with different modules perform on a node-wise prediction. Nodes were selected all across the mesh, to provide a comprehensive overview of the prediction capabilities on different zones of the mesh. Results correspondent to selected nodes all over the aneurysm mesh are reported in Figures 7-11.
@@ 3.4 Extrapolation task: numerical results
For the extrapolation task we show the prediction on the normalized WSI (at systolic peak) and OSI (at the final timestamp) with and without thresholds (Figures 20, 21 and 22, respectively).
@@ 4 Conclusion
In this paper we developed a novel framework for the prediction of shear stress indices, such as the WSS, integrating Full Order Models techniques with surrogate models; in particular we obtained high fidelity data from blood flow modeling via Finite Volumes discretization in order to train Graph Neural Networks. This approach overcomes the traditional mesh refinement problem, as GNNs can yield instantaneous evaluations on the input meshes, which can be made of different numbers of nodes. GNNs exhibit good qualitative performance in predicting WSS and OSI, both on interpolation and extrapolation tasks, providing an alternative tool to monitor shear stress peaks at different volume sizes and consequently to quantify the risk of ruptures or dissections in the TAA.
Our method is not free from open issues. Firstly, the implemented GNNs are permutation equivariant, but not SE3 equivariant. Although our results don’t show limitations arising from the lack of this property, the implementation of SE3 equivariance would make our approach more robust to data rotation (see, for instance, the approach used in [48]). The individual performances of the nodes need to be enhanced to ensure accurate local predictions. One way to achieve this is by treating the data as temporal sequences and using temporal models, namely, Temporal GNNs [49]. This approach would involve a significant preprocessing stage, where we’d need to define a sliding window of a certain length to create temporal graph sequences that are long enough to capture temporal behavior, yet short enough to be processed by a relatively simple Temporal GNN, in order to avoid issues like vanishing gradients [50]. Another option would be to process the high-resolution data as point clouds, allowing us to apply state-of-the-art methods for prediction tasks on point clouds, such as PointNets [51]. Eventually, with more advanced hardware sources, we could fine-tune our model via grid-search methods to find the optimal configuration. Another notable limitation of this study is the reliance on data from a single patient. While the results provide valuable insights, they may not fully capture the variability across a broader population. Future work should focus on validating the findings using data from multiple patients to enhance the generalizability of the proposed approach. Expanding the dataset will also allow for a more robust assessment of inter-patient variability and its implications on the model’s predictive performance.
@@ Acknowledgements
G.A.D., P.C.A., and G.R. acknowledge the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. S.S. and G.R. acknowledge the support provided by PRIN ”FaReX - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning” project, and by INdAM-GNCS 2019–2020 projects and PON ”Research and Innovation on Green related issues” FSE REACT-EU 2021 project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them. In addition, the authors would like to acknowledge INdAM–GNCS.

### v3

@@ Abstract
The complexity of the cardiovascular system needs to be accurately reproduced in order to promptly acknowledge health conditions; to this aim, advanced multifidelity and multiphysics numerical models are crucial. On one side, Full Order Models (FOMs) deliver accurate hemodynamic assessments, but their high computational demands hinder their real-time clinical application. In contrast, Reduced Order Models (ROMs) provide more efficient yet accurate solutions, essential for personalized healthcare and timely clinical decision-making. In this work, we explore the application of computational fluid dynamics (CFD) in cardiovascular medicine by integrating FOMs with ROMs for predicting the risk of aortic aneurysm growth and rupture. Wall Shear Stress (WSS) and the Oscillatory Shear Index (OSI), sampled at different growth stages of the thoracic aortic aneurysm, are predicted by means of Graph Neural Networks (GNNs). GNNs exploit the natural graph structure of the mesh obtained by the Finite Volume (FV) discretization, taking into account the spatial local information, regardless of the dimension of the input graph. Our experimental validation framework yields promising results, confirming our method as a valid alternative that overcomes the curse of dimensionality.
@@ 1 Introduction
Thoracic aortic aneurysms (TAAs) represent a severe form of aneurysm that manifests as unrepairable bulges in the thoracic aorta. These are particularly prevalent among the elderly, necessitating careful and regular monitoring of their size, especially in the case of small TAAs [3]. The progression and potential rupture of aneurysms are critical concerns, and their management often involves balancing the risks of surgical intervention against the potential for catastrophic vascular failure.
@@ 2.1 FOM
This section introduces the mathematical model used to describe blood flow in the toracic aorta, specifically the incompressible Navier-Stokes equations. This includes adopted boundary conditions and discretization methods.
@@ 2.1.1 Mathematical Model
The blood motion is determined within a time-independent domain $\Omega$ over a time interval $[0,T]$. We assume, without loss of generality, that the domain boundary $\partial\Omega$ is partitioned into 4 subsets as $\partial\Omega=\Gamma^{\mathrm{in}}\cup\Gamma^{\mathrm{out}}_{\mathrm{SA}}\cup\Gamma^{\mathrm{out}}_{\mathrm{abd}}\cup\Gamma^{\mathrm{wall}}$, as shown in Figure 2. The flow is governed by the following equations:
$\begin{array}[]{ll}\rho\partial_{t}\boldsymbol{u}+\rho\nabla\cdot(\boldsymbol{u}\otimes\boldsymbol{u})-\nabla\cdot\boldsymbol{\sigma}=0&\text{ in }\Omega\times\left(0,T\right],\end{array}$ (1)
@@
where $\rho=1060\,\text{kg/m}^{3}$ is the blood density, $\mathbf{u}$ represents the velocity vector, and $\boldsymbol{\sigma}$ denotes the Cauchy stress tensor. Eq. 1 expresses the conservation of momentum, while Eq. 2 ensures the mass conservation. The blood is modeled as a non-Newtonian fluid, where $\boldsymbol{\sigma}$ is defined as:
@@
where $\mathbf{D}=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{\mathrm{T}}\right)$, $p$ represents the pressure, $\mu$ is the dynamic viscosity, and $\dot{\gamma}$ is the shear rate:
@@
where $\boldsymbol{\tau}_{0}$ and $\mu_{\infty}$ are 0.005 and 0.0035, respectively. Accounting for the non-Newtonian rheology of blood enables a more precise characterization of the hemodynamic indicators described below in the thoracic aorta, particularly under pathological conditions, as demonstrated in [18, 19].
@@
where $\mathbf{n}$ is the normal vector, and $\boldsymbol{\tau}_{w}:=\boldsymbol{\tau}_{|\Gamma^{\mathrm{wall}}}$ denotes the wall shear stress tensor.
@@
For the sake of simplicity, in what follows we denote by WSS the magnitude of $\mathbf{WSS}$.
@@ 2.1.3 Numerical discretization
To evaluate the independence of CFD results from mesh size, four sets of meshes containing $1.2\times 10^{5}$, $2.4\times 10^{5}$, $3.3\times 10^{5}$ and $4.5\times 10^{5}$ tetrahedral elements were created. These meshes correspond to varying element sizes on the aneurysm wall, specifically $0.3$ mm, $0.2$ mm, $0.15$ mm, and $0.125$ mm, respectively. Temporal variations of the area-averaged WSS on the aneurysm wall for all mesh sizes are presented in Fig. 3. This figure shows minimal differences between the two finest meshes, leading to the selection of the third mesh for subsequent simulations.
@@
Three cardiac cycles are required to achieve fully periodic flow. Additionally, to ensure convergence, initial values of zero for pressure and velocity were applied, making the early cycles inaccurate.
@@ 3.1 Experimental setup
    Connectivity: The connectivity of the surface mesh of the TAA is extracted from the Finite Volume Discretization, leading to the reference graph. The number of nodes and the connectivity depends on the specific size; values are reported in Table 1 ;
@@
    Node features: for each timestamp and for each TAA size, we label the graph nodes $v$ with a feature $\mathbf{x}_{v}=[x,y,z,t,u(t),p_{\text{abd}}(t),p_{\text{SA}}(t)]$, where: $x,y,z$ are the coordinates of each node (that varies according to the TAA size), $t$ is the times step, $u(t)$ is the inlet flow rate, $p_{\text{abd}}(t),p_{\text{SA}}(t)$ are the abdominal and SA outlet pressure.
@@
    the first step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 100%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 75%. This task focuses on the interpolation capabilities of the model over shear indices at different growth stages of the TAA.
@@
    the second step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 75%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 100%. This task focuses on the extrapolation capabilities of the model over shear indices at different growth stages of the TAA. This task is clinically more relevant as we would like to predict the WSS on future (and therefore larger) volume sizes of the TAA.
@@
From a qualitative point of view, we compare how GNNs with different modules perform on a node-wise prediction. Nodes were selected all across the surface mesh, to provide a comprehensive overview of the prediction capabilities on different zones of the aneurysm wall. Results correspondent to selected nodes all over the aneurysm surface mesh are reported in Figures 7-11.
@@ 3.4 Extrapolation task: numerical results
For the extrapolation task we show the prediction on WSS (at systolic peak) and OSI (at the final timestamp) with and without thresholds (Figures 20, 21 and 22, respectively). Both indices are normalized in $[0,1]$.
@@ 4 Conclusion
In this paper, we developed a novel framework for the prediction of shear stress indices, such as the WSS, integrating Full Order Models techniques with surrogate models; in particular, we obtained high fidelity data from blood flow modeling via Finite Volumes discretization in order to train Graph Neural Networks. This approach overcomes the traditional mesh refinement problem, as GNNs can yield instantaneous evaluations on the input meshes, which can be made of different numbers of nodes. GNNs exhibit good qualitative performance in predicting WSS and OSI, both on interpolation and extrapolation tasks, providing an alternative tool to monitor shear stress peaks at different volume sizes and consequently to quantify the risk of ruptures or dissections in the TAA.
Our method is not free from open issues. Firstly, the implemented GNNs are permutation equivariant, but not SE3 equivariant. Although our results do not show limitations arising from the lack of this property, the implementation of SE3 equivariance would make our approach more robust to data assimilation (see, for instance, the approach used in [50]). As our model leverages space coordinates as relevant features for an effective prediction, same data taken in a different system of coordinates could be subject to a bad predictive behaviour. A SE3 equivariant model would inherently possess the desirable feature of being equivariant to rototranslations in space.
The individual performances of the nodes need to be enhanced to ensure accurate local predictions. One way to achieve this is by treating the data as temporal sequences and using temporal models, namely, Temporal GNNs [51]. This approach would involve a significant preprocessing stage, where we’d need to define a sliding window of a certain length to create temporal graph sequences that are long enough to capture temporal behavior, yet short enough to be processed by a relatively simple Temporal GNN, in order to avoid issues like vanishing gradients [52]. Another option would be to process the high-resolution data as point clouds, allowing us to apply state-of-the-art methods for prediction tasks on point clouds, such as PointNets [53]. Eventually, with more advanced hardware sources, we could fine-tune our model via grid-search methods to find the optimal configuration. Another notable limitation of this study is the reliance on data from a single patient. While the results provide valuable insights, they may not fully capture the variability across a broader population. Future work should focus on validating the findings using data from multiple patients to enhance the generalizability of the proposed approach. Expanding the dataset will also allow for a more robust assessment of inter-patient variability and its implications on the model’s predictive performance.
@@ Acknowledgements
G.A.D., P.C.A., and G.R. acknowledge the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project and its CC5 Young Researchers initiative. S.S. and G.R. acknowledge the support provided by PRIN ”FaReX - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning” project, and by INdAM-GNCS 2019–2020 projects and PON ”Research and Innovation on Green related issues” FSE REACT-EU 2021 project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them. P.C.A. also acknowledges the INdAM-GNCS Project “Metodi numerici efficienti per problemi accoppiati in sistemi complessi” (CUP E53C24001950001). In addition, the authors would like to acknowledge INdAM–GNCS.